# Competition between transient oscillations and early stochasticity in exponentially growing populations

### v2

@@ VII Discussions and future directions
On the theoretical side, the competition between the transient dynamics and early stochasticity is what we believe to be a much more general phenomenon in stochastic exponential growth models. In most models of stochastic exponential growth, the early phases of growth are highly stochastic, but as the size of the system grows, the dynamics becomes more deterministic. In such cases, the effect of early stochasticity freezes in the distribution of the exponentially growing state. A well-known example of such dynamics is the Yule process, where the agents can spontaneously divide into two at a constant rate in a Markovian fashion. While the early stages of the Yule process are highly stochastic, it asymptotically leads to deterministic exponential growth with a pre-factor that approaches a stationary distribution. Similarly, Branching processes such as the Bellman-Harris model of population growth [38] where the generation times are independently drawn from a fixed distribution, show early stochasticity that freezes in the final state of the system. Stochastic exponential growth can be also studied in the context of stochastic differential equations (SDEs) such as the Feller square root process [45] and its generalizations to power-law noise [46]. Similar effects can be observed in these models as long as the stochasticity grows slower than the system size and they can be mapped to square root noise through a change of variable [47]. The exception is the case of geometric Brownian motion, where the noise grows proportional to the system’s size [48]. Stochastic exponential growth is also observed in autocatalytic reaction networks such as those used in models of the origin of life [49, 50] and models of single-cell growth [51, 52]. These models also have early stochasticity that grows asymptotically with the square root of the system size.
@@
Besides in exponentially growing systems, both synchronized divisions and finite-size stochasticity are present in theoretical models and experimental setups of systems with finite population sizes such as the Moran process and turbidostat. It has recently been shown that the oscillations in the division rate in such systems persist indefinitely due to the finite coalescent time of finite populations [53]. While the transition is not present in these systems due to the lack of exponential growth, we hope the techniques used in this paper prove fruitful in the analysis of such systems. Finite-size models are important in the context of population genetics since populations cannot grow exponentially over evolutionary timescales. Understanding the interaction of transient dynamics controlled by single-cell statistics and stochasticity in finite-size populations can provide insight into how evolution acts on cellular physiology.
@@ Appendix B The transient dynamics of total mass and synchronicity
for some constant $A_{\text{ens}}$, where $r=\Lambda_{\infty}-\text{Re}\mu$ and $\Omega=\text{Im}\mu$. Plugging this into the definition of $\Lambda_{\text{ens}}(t)$ from (6) yields the first-order behavior of Eq. (7) after ignoring terms of order $O(e^{-2rt})$.
@@
where $\langle$.$\rangle$ . ⟩ is essentially a path integral over growth state space that integrates $\lambda_{t}$ over all growth rate trajectories along a lineage. The leading complex-valued constant $\mu$ for which asymptotically $\Phi(t)\rangle\propto e^{\mu t}$ can be obtained from this expression via
@@
Appendix C The oscillation amplitude
In this section, we will discuss how the oscillation amplitude prefactor in Eq. (12) is distributed and derive its typical value. We will assume that $\lambda_{t}$ is an Ornstein-Uhlenbeck process with mean $\bar{\lambda}$, variance $\sigma_{\lambda}^{2}$ and correlation time $\tau_{\text{cor}}$. To find the amplitude of oscillations in cell count growth rate $\Lambda(t)$, we assume the population is large enough to the point where $\partial_{t}M(t)=\Lambda_{\infty}M(t)$ and $\partial_{t}\Phi(t)=\mu\Phi(t)$. We then plug in Eq. (22) into the definition of $\Lambda(t)$ to find
@@ Appendix C The oscillation amplitude
$\Lambda(t)\approx\Lambda_{\infty}\left(1+2\text{Re}\left[e^{(i\omega-\frac{1}{2}\omega^{2})\text{CV}_{m_{d}}^{2}}\Psi(t)\right]\right),$ (33)
where we made an approximation $(\mu/\Lambda_{\infty}-1)/(1+i\omega)\approx 1$ for simplicity. Recall that the cell mass synchronicity is given by $\Psi(t)=\Phi(t)/M(t)$. Eq. (33) tells us that the instantaneous amplitude is equal to $2e^{-\frac{1}{2}\omega^{2}\text{CV}^{2}_{m_{d}}}|\Psi(t)|$. This is how we determined the instantaneous amplitude trajectories in Fig. 3 (a) and (c). The oscillation fits in Fig. 2 and Fig. 1 were obtained from calculating the simulated cell mass synchronicity $\Psi(t_{max})$ at the final simulation time $t_{max}=19\tau_{\text{div}}$ and extrapolating its expected trajectory back in time by plugging $\Psi(t)\approx\Psi(t_{\text{max}})e^{(-r+i\Omega)(t-t_{\text{max})}}$ into Eq. (33). In this procedure, as $t_{\text{max}}$ increases, one can obtain better estimates of the asymptotic amplitude pre-factor $A$ and phase $\phi$. In fact, by matching Eq. (33) to Eq. (12), we can obtain a formal definition of the random asymptotic amplitude pre-factor and phase as a function of $\Psi(t)$ at infinity,
$Ae^{i\phi}:=\lim_{t\to\infty}e^{(i\omega-\frac{1}{2}\omega^{2})\text{CV}_{m_{d}}^{2}}e^{(r-i\Omega)t}\Psi(t).$ (34)
We will use this definition to analyze the properties of $A$ as a random variable. similarly to Eq. (34), one can show that the ensemble average population oscillation amplitude and phase from Eq. (7) satisfy
$A_{\text{ens}}e^{i\phi_{ens}}=\lim_{t\to\infty}e^{(i\omega-\frac{1}{2}\omega^{2})\text{CV}_{m_{d}}^{2}}e^{(r-i\Omega)t}\frac{\langle\Phi(t)\rangle}{\langle M(t)\rangle}.$ (35)
@@
$A_{\text{ens}}=2e^{-\frac{1}{2}\omega^{2}\text{CV}_{m_{d}}^{2}+\frac{3}{2}\omega^{2}\sigma_{\lambda}^{2}\tau_{\text{cor}}^{2}}|\Psi(0)|.$ (38)
The term $\frac{3}{2}\omega^{2}\sigma_{\lambda}^{2}\tau_{\text{cor}}^{2}$ results from the constant shift in Eq. (37).
We are now interested in the distribution and typical value of $A$ Let us assume that variability in $M(t)$ is negligible. We can write
$Ae^{i\phi}\approx\lim_{t\to\infty}e^{(i\omega-\frac{1}{2}\omega^{2})\text{CV}_{m_{d}}^{2}}e^{-\mu t}\frac{\Phi(t)}{M(0)}.$ (39)
Note that $\langle Ae^{i\phi}\rangle=A_{\text{ens}}e^{i\phi_{\text{ens}}}$. This already gives us the first indication that the average amplitude of a single simulation is strictly higher than the ensemble average amplitude, by noting that
$\langle A\rangle=\langle|Ae^{i\phi}|\rangle>|\langle Ae^{i\phi}\rangle|=A_{\text{ens}}.$ (40)
@@
which is a complex-valued random variable with $\langle Z(m_{0})\rangle=1$. The first division takes place at $t=T_{m_{0}}$, at which the population splits up into two identical sub-populations that start with mass $1$ at time $t=T_{m_{0}}$. This gives us a relationship for the collective mass phase similar to the one used to derive Powell’s relationship [9].
@@
where the arrows are used to distinguish the two sub-populations. Plugging this relationship into Eq. (44) we find
@@
which gives the typical amplitude for a population stemming from one cell. When the population starts from multiple cells with masses $m_{j}$, the collective mass phase is a complex superposition of the mass phase trajectories of the populations stemming from each of the individual cells $\Phi_{j}(m_{j},t)$, so
@@
We can now plug this into Eq. (39) and use the single-cell result to obtain
$\bar{A}=2e^{-\frac{1}{2}\omega^{2}\text{CV}_{m_{b}}^{2}+3\omega^{2}\sigma_{\lambda}^{2}\tau_{\text{cor}}^{2}}\sqrt{|\Psi(0)|^{2}+J(0)},$ (56)
@@
$J(0)=\frac{1}{M_{0}^{2}}\sum_{j=1}^{N(0)}m_{j}^{2}\left(\frac{m_{j}^{-\epsilon}}{2^{1-\epsilon}-1}-1\right).$ (57)
For all values of $\epsilon$ and $m_{j}$, this is well approximated by Eq. (20) in the main text. When the initial population has a well-mixed mass distribution and $\Phi_{\text{total}}(0)$ is close to zero, one could argue with the help of a central limit theorem that $\Phi_{\text{total}}(t)$ and $Ae^{i\phi}$ have a centered rotationally symmetric bi-variate Gaussian distribution in the complex plane. In that case, we know that $A$ itself must have a $\chi_{2}$-distribution, with a probability density function of
@@
Table 2: Experimentally measured values of the mean, standard deviation, and mother-daughter correlations of per cell growth rates for E. coli in several growth media as reported in Table S3 of Ref. [27]. The mother-daughter correlations were obtained from the elongation rates in Figure S4 of Ref. [27] by taking the first-generation correlations and subtracting the average of the three final correlations as extrinsic noise.
Appendix D Determining the growth process parameters
In Ref. [25], a relationship is derived between parameters of an Ornstein-Uhlenbeck growth process $\bar{\lambda}$, $\sigma_{\lambda}$, $\tau_{\text{cor}}$ and the emergent mean, $\bar{\kappa}$, standard-deviation $\sigma_{\kappa}$ and mother-daughter correlations $\rho_{m-d}$ of growth rate averaged over cell cycles. Up to first order in $\sigma_{\lambda}^{2}/\bar{\lambda}^{2}$, their results are
@@ Appendix D Determining the growth process parameters
By inverting this set of equations we obtained the Ornstein-Uhlenbeck process parameters in Table 1 based on values from Ref. [27]. The values used are listed in Table 2.
@@
To show that our results are consistent with the transition point derived in the introduction, we want to take a limit of the continuous growth rate model that corresponds to the per-cell growth rate model with vanishing correlations from Ref. [1]. This can be achieved by taking $\bar{\lambda}\tau_{\text{cor}}\ll 1$ while fixing $D=\sigma_{\lambda}^{2}\tau_{\text{cor}}/\bar{\lambda}$. In this limit we obtain

### v3

@@ VII Discussions and future directions
On the theoretical side, the competition between the transient dynamics and early stochasticity is what we believe to be a much more general phenomenon in stochastic exponential growth models. In most models of stochastic exponential growth, the early phases of growth are highly stochastic, but as the size of the system grows, the dynamics becomes more deterministic. In such cases, the effect of early stochasticity freezes in the distribution of the exponentially growing state. A well-known example of such dynamics is the Yule process, where the agents can spontaneously divide into two at a constant rate in a Markovian fashion. While the early stages of the Yule process are highly stochastic, it asymptotically leads to deterministic exponential growth with a pre-factor that approaches a stationary distribution. This pre-factor is the aforementioned frozen effect of stochasticity in the final state of the system. Similarly, Branching processes such as the Bellman-Harris model of population growth Bellman and Harris (1952) where the generation times are independently drawn from a fixed distribution, show early stochasticity that freezes in the final state of the system. Stochastic exponential growth can be also studied in the context of stochastic differential equations (SDEs) such as the Feller square root process Feller (1951) and its generalizations to power-law noise Cox (1996). Similar effects can be observed in these models as long as the stochasticity grows slower than the system size and they can be mapped to square root noise through a change of variable Pirjol et al. (2017). The exception is the case of geometric Brownian motion, where the noise grows proportional to the system’s size Van Kampen (1992). Stochastic exponential growth is also observed in autocatalytic reaction networks such as those used in models of the origin of life Jafarpour et al. (2017, 2015) and models of single-cell growth Iyer-Biswas et al. (2014); Lin et al. (2020). These models also have early stochasticity that grows asymptotically with the square root of the system size.
@@
Besides in exponentially growing systems, both synchronized divisions and finite-size stochasticity are present in theoretical models and experimental setups of systems with finite population sizes such as the Moran process and turbidostat. It has recently been shown that the oscillations in the division rate in such systems persist indefinitely due to the finite coalescent time of finite populations Jafarpour et al. (2023). While the transition is not present in these systems due to the lack of exponential growth, we hope the techniques used in this paper prove fruitful in the analysis of such systems. Finite-size models are important in the context of population genetics since populations cannot grow exponentially over evolutionary timescales. Understanding the interaction of transient dynamics controlled by single-cell statistics and stochasticity in finite-size populations can provide insight into how evolution acts on cellular physiology.
In many natural settings, bacteria grow in colonies with spatial structures. Within a spatially structured colony, cells that are closely related in their population tree are also closely related in space. This could create a strong local synchronization, despite the spatial heterogeneity that can desynchronize cells that are spatially separated. The interplay between this local synchrony and the mechanical and biological properties of the colony is a subject of future research.
@@ Appendix B The transient dynamics of total mass and synchronicity
for some constants $A_{\text{N,ens}}$ and $\phi_{\text{N,ens}}$, where $r=\Lambda_{\infty}-\text{Re}\mu$ and $\Omega=\text{Im}\mu$. Plugging this into the definition of $\Lambda_{\text{ens}}(t)$ from (6) yields the first-order behavior of Eq. (7) after ignoring terms of order $O(e^{-2rt})$.
@@
where $\langle$.$\rangle$ . ⟩ is essentially a path integral over growth state space that integrates $\lambda_{t}$ over all growth rate trajectories along a lineage. The leading complex-valued constant $\mu$ for which asymptotically $\langle\Phi(t)\rangle\propto e^{\mu t}$ can be obtained from this expression via
@@
Appendix C The binned population growth rate
In this section, we discuss how we estimated the population growth rate, as well as oscillation fits and fluctuation envelopes shown in Fig. 1 and Fig. 2. Naturally, when dealing with finite discrete populations, the population growth rate in Eq. (10) is not well defined, since it contains the derivative of a discrete variable. We thus consider small but finite time intervals $\Delta t$. The discretized population growth rate shown in Fig. 1 and Fig. 2 is defined as
$\Lambda_{\Delta t}=\frac{\ln(N(t+\Delta t/2))-\ln(N(t-\Delta t/2))}{\Delta t}$ (33)
C.1 The fluctuation envelope
In the limit of a population completely dominated by random fluctuations, all collective synchronicity disappears and we may assume the cells in the population to be independent. Consider a population of $N(t-\Delta t/2)$ independent cells. Given a small enough time interval $\Delta t$, each cell randomly chooses to divide or not divide with the same probability $p$. The number of cells after this time interval $N(t+\Delta t/2)$ can now be related to the previous population via
$N(t+\Delta t/2)=N(t-\Delta t/2)+\mathcal{B}$ (34)
where $\mathcal{B}\sim\text{Binomial}(N(t-\Delta t/2),p)$. Assuming that the expected population predictably grows as $\langle N(t+\Delta t/2)\rangle=e^{\Lambda\infty\Delta t}\langle N(t-\Delta t/2)\rangle$, we can derive the probability that an individual cell divides as $p\approx\Lambda_{\infty}\Delta t$. The discretized population growth rate can be expressed as
$\Lambda_{\Delta t}=\frac{1}{\Delta t}\ln\left(1+\frac{\mathcal{B}}{N(t-\Delta t/2)}\right)$ (35)
By using properties of the binomial distribution, one can show that up to leading order in small $\Delta t$, typical fluctuations in $\Lambda_{\Delta t}$ scale with
@@ C.1 The fluctuation envelope
$\sqrt{\text{Var}\left(\Lambda_{\Delta t}\right)}\approx\sqrt{\frac{\Lambda_{\infty}}{N(t)\Delta t}}.$ (36)
To smoothen out this fluctuation envelope, we approximated the cell count by the scaled total cell mass $N(t)\approx M(t)/2\ln(2)$, to obtain
$\sqrt{\text{Var}\left(\Lambda_{\Delta t}\right)}\approx\sqrt{\frac{\Lambda_{\infty}2\ln(2)}{M(t)\Delta t}}.$ (37)
This is what we used to determine the fluctuation envelope in Fig. 1 (c) and (d).
C.2 The oscillation fit and envelope
In this section, we explain how we determined the oscillation fits and envelopes Fig. 1 and Fig. 2. Usually, a fit would involve a procedure that minimizes the distance between data and a parametrized curve. Our method instead employs the idea that oscillation phase and amplitude can be directly deduced from the population cell mass distribution at any point in time. Since the oscillations become more and more deterministic as the population gets larger, the best estimate of the transient limit amplitude and phase are based on the mass distribution at the largest simulated time $t_{max}$ available.
In the limit of a large population, the total cell mass and collective mass phase will approach the dynamics of their expectation values, hence for large $t$ we have $\partial_{t}M(t)=\Lambda_{\infty}M(t)$ and $\partial_{t}\Phi(t)=\mu\Phi(t)$. For all large $t$ we can reasonably extrapolate these deterministic dynamics around the population’s state at some large time $t_{max}$ to obtain
$M(t)\approx M(t_{max})e^{\Lambda_{\infty}(t-t_{max})},$ (38)
and
$\Phi(t)\approx\Phi(t_{max})e^{\mu(t-t_{max})}.$ (39)
Subsequently, we can plug these approximations into Eq. (33) using a first-order expansion of Eq. (22) to find
$\Lambda_{\Delta t}\approx\Lambda_{\infty}\left(1+\text{Re}\left[C_{\Delta t}\Psi(t_{max})e^{(-r+i\Omega)(t-t_{max})}\right]\right),$ (40)
where $\Psi(t)=\Phi(t)/M(t)$ and
$C_{\Delta t}=2e^{(i\omega-\frac{1}{2}\omega^{2})\text{CV}_{m_{d}}^{2}}\frac{e^{(-r+i\Omega)\Delta t/2}-e^{-(-r+i\Omega)\Delta t/2}}{\Delta t(1+i\omega)\Lambda_{\infty}},$ (41)
in the main text, we use the approximation
$C_{\Delta t}\approx 2e^{(i\omega-\frac{1}{2}\omega^{2})\text{CV}_{m_{d}}^{2}},$ (42)
which is valid for small $D$ and $\Delta t$. Determining the envelope of the oscillations in Eq. (40) is straightforward: You take the absolute value of the term that rotates in the complex plane. The upper envelope is thus given by.
$\Lambda_{env}\approx\Lambda_{\infty}\left(1+|C_{\Delta t}\Psi(t_{max})|e^{-r(t-t_{max})}\right)$ (43)
Subsequently, we define the instantaneous amplitude at time $t$ as
$A_{\text{inst}}(t)=|C_{\Delta t}\Psi(t)|$ (44)
Given a population at time $t_{max}$ with dynamics that are like a large population, the best estimate for the amplitude of future oscillations is $A_{\text{inst}}(t_{max})e^{-r(t-t_{max})}$.
Appendix D The amplitude distribution
In the previous section, we discussed how we can fit transient oscillations to the population growth rate based on the instantaneous complex amplitude $C_{\Delta t}\Psi(t_{max})$ at some large time $t_{max}$. As we increase $t_{max}$, Eq. (40) becomes an increasingly better descriptor of the transient oscillations at large $t$. This is demonstrated in Fig. 3 (a) and (c), where we show trajectories of the instantaneous amplitudes $A_{\text{inst}}(t)=|C_{\Delta t}\Psi(t)|$ of independent populations at finite times. Importantly, we see that at large $t$, all trajectories approach a curve of the form $Ae^{-rt}$. Similarly, the complex instantaneous amplitude $C_{\Delta t}\Psi(t)$ approaches a complex curve of the form $Ae^{i\phi}e^{-(r-i\Omega)t}$. Consequently, Eq. (40) can be written as
$\Lambda(t)\approx\left(1+Ae^{-rt}\cos(\Omega t+\phi)\right),$ (45)
where amplitude pre-factor $A$ and phase $\phi$ can be expressed as the limit
$Ae^{i\phi}=\lim_{t\to\infty}C_{\Delta t}\Psi(t)e^{(r-i\Omega)t}.$ (46)
In this section, we will use this limit to determine the typical transient amplitude $\bar{A}=\sqrt{\langle A^{2}\rangle}$ and its distribution in a special case where initial cell masses are out of phase. In Eq. (46) we see that $Ae^{i\phi}$ is proportional to $\Psi(t)=\Phi(t)/M(t)$ at large times $t$, where we recall that $\Phi(t)=\sum_{j}m_{j}^{1+i\omega}$ is the collective mass phase and $M(t)=\sum_{j}m_{j}$ is the total cell mass. Variability in $M(t)$ is very small compared to the variability in $\Phi(t)$, so we will approximate the total cell mass by its expected value $\Psi(t)\approx\Phi(t)/\langle M(t)\rangle$. Our analysis of $A$ will thus use the expression
$Ae^{i\phi}\approx\lim_{t\to\infty}C_{\Delta t}\frac{\Phi(t)}{\langle M(t)\rangle}e^{(r-i\Omega)t}$ (47)
$A$ and $\phi$ are closely related to the amplitude $A_{\text{ens}}$ and phase $\phi_{\text{ens}}$ of the ensemble population.
By taking the expectation value on both sides of Eq. (22) and plugging the first-order expansion into Eq. (6), we obtain first-order oscillatory behavior of the form Eq. (7), where the ensemble amplitude $A_{\text{ens}}$ and phase $\phi_{\text{ens}}$ are given by
$A_{\text{ens}}e^{i\phi_{\text{ens}}}=\lim_{t\to\infty}C_{\Delta t}\frac{\langle\Phi(t)\rangle}{\langle M(t)\rangle}e^{(r-i\Omega)t}$ (48)
@@ Appendix D The amplitude distribution
$A_{\text{ens}}=e^{\frac{3}{2}\omega^{2}\sigma_{\lambda}^{2}\tau_{\text{cor}}^{2}}|C_{\Delta t}\Psi(0)|.$ (51)
The term $\frac{3}{2}\omega^{2}\sigma_{\lambda}^{2}\tau_{\text{cor}}^{2}$ results from the constant shift in Eq. (50). From Eq. (47) and Eq. (48) we can tell that $\langle Ae^{i\phi}\rangle\approx A_{\text{ens}}e^{i\phi_{\text{ens}}}$. This already gives us the first indication that the average amplitude of a single simulation is strictly higher than the ensemble average amplitude, by noting that
$\langle A\rangle=\langle|Ae^{i\phi}|\rangle>|\langle Ae^{i\phi}\rangle|\approx A_{\text{ens}}.$ (52)
@@
which is a complex-valued random variable with $\langle Z(m_{0})\rangle=1$. We use notation $Z=Z(1)$. The first division takes place at $t=T_{m_{0}}$, at which point the population splits up into two identical sub-populations that start with mass $1$ at time $t=T_{m_{0}}$. This gives us a relationship for the collective mass phase similar to the one used to derive Powell’s relationship Powell (1956).
@@
where the arrows are used to distinguish the two independent sub-populations. Plugging this relationship into Eq. (56) we find
@@
which gives the typical amplitude for a population stemming from one cell.
D.1 Multiple cells in initial population
When the population starts from multiple cells with masses $m_{j}$, the collective mass phase is a complex superposition of the mass phase trajectories of the populations stemming from each of the individual cells $\Phi_{j}(m_{j},t)$, so
@@ D.1 Multiple cells in initial population
Note that the different mass phase trajectories $\Phi_{j}(m_{j},t)$ are independent. We can thus plug this into (47) and take the expected absolute square to find
$\bar{A}=|C_{\Delta t}|e^{\frac{3}{2}\omega^{2}\sigma_{\lambda}^{2}\tau_{\text{cor}}^{2}}\sqrt{|\Psi(0)|^{2}+J(0)},$ (68)
@@
$J(0)=\frac{1}{M(0)^{2}}\sum_{j=1}^{N(0)}m_{j}^{2}\left(\frac{m_{j}^{-\epsilon}}{2^{1-\epsilon}-1}-1\right).$ (69)
For all values of $\epsilon$ and $m_{j}$, this is well approximated by Eq. (20) in the main text. We also approximated $C_{\Delta t}$ by Eq. (42). When the initial population has a well-mixed mass distribution and $\Phi_{\text{total}}(0)$ is close to zero, one could argue with the help of a central limit theorem that $\Phi_{\text{total}}(t)$ and $Ae^{i\phi}$ have a centered rotationally symmetric bi-variate Gaussian distribution in the complex plane. In that case, we know that $A$ itself must have a $\chi_{2}$-distribution, with a probability density function of
@@
Table 2: Experimentally measured values of the mean, standard deviation, and mother-daughter correlations of per cell growth rates for E. coli in several growth media as reported in Table S3 of Ref. Taheri-Araghi et al. (2015). The mother-daughter correlations were obtained from the elongation rates in Figure S4 of Ref. Taheri-Araghi et al. (2015) by taking the first-generation correlations and subtracting the average of the three final correlations as extrinsic noise.
Appendix E Determining the growth process parameters
In Ref. Hein and Jafarpour (2022), a relationship is derived between parameters of an Ornstein-Uhlenbeck growth process $\bar{\lambda}$, $\sigma_{\lambda}$, $\tau_{\text{cor}}$ and the emergent mean, $\bar{\kappa}$, standard-deviation $\sigma_{\kappa}$ and mother-daughter correlations $\rho_{m-d}$ of growth rate averaged over cell cycles. Up to first order in $\sigma_{\lambda}^{2}/\bar{\lambda}^{2}$, their results are
@@ Appendix E Determining the growth process parameters
By inverting this set of equations we obtained the Ornstein-Uhlenbeck process parameters in Table 1 based on values from Ref. Taheri-Araghi et al. (2015). The values used are listed in Table 2.
@@
To show that our results are consistent with the transition point derived in the introduction, we want to take a limit of the continuous growth rate model that corresponds to the per-cell growth rate model with vanishing correlations from Ref. Jafarpour (2019). This can be achieved by taking $\bar{\lambda}\tau_{\text{cor}}\ll 1$ while fixing $D=\sigma_{\lambda}^{2}\tau_{\text{cor}}/\bar{\lambda}$. In this limit we obtain